40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7870     Entry time: Fri Dec 21 19:49:39 2012
Author: Riju 
Type: Update 
Category:  
Subject: Photodiode transimpedance 

I have repeated the transimpedance measurement of PDA10CF. Also made the dark current noise measurement by connecting the PDA10CF output to the A channel of network analyzer.  The results are as follows. I I started to take the reading for shot noise intercept current using a light bulb in front of the PD, changing the current through the bulb, but at higher current the bulb filament got broken, so the experiment is incomplete.

Attachment 1: PDA10CFrepeat.pdf  9 kB  | Hide | Hide all
PDA10CFrepeat.pdf
Attachment 2: darknoiseVpda10cf.pdf  4 kB  | Hide | Hide all
darknoiseVpda10cf.pdf
Attachment 3: darknoiseApda10cf.pdf  4 kB  | Hide | Hide all
darknoiseApda10cf.pdf
Attachment 4: PDA10CF_z.pdf  4 kB  | Hide | Hide all
PDA10CF_z.pdf
ELOG V3.1.3-